if f(x) =x/3 +5 what is the value of f^-1 (-6)
Answer by stanbon(75887)

if f(x) =x/3 +5 what is the
value of f^-1 (-6)
----
Let y = -6 and solve for "x":
-6 = (x/3)+5
x/3 = -11
x = -33
------------------------------
f^-1(-6) = -33
==================
Cheers,
Stan H.
=============